# Robust Utility Optimization via a GAN Approach

### v1

@@ Results
NN based strategies are able to adapt to concrete real-world settings and learn to invest more in risky assets, only if drift and diffusion characteristics allow for more potential upside over the dedicated trading time horizon, compared to the risk-less bank account investment. Depending on the time horizon, $\pi^{\text{NN}}_{\text{non rob}}$ might look very different even if annual drift and diffusion characteristics are held at the same level. The asymptotic strategy $\pi_{\text{ref}}$ however, keeps proportional investments in the no-trade region corresponding to those annual characteristics, irrespective of the time horizon. Therefore we suspect that for reasonable configurations of market characteristics, the difference between the NN-based $\pi^{\text{NN}}_{\text{non rob}}$ and the asymptotic solution $\pi^{\text{ref}}$ is most prominent in short-term trading.
More concretely, the risky asset $S$ has an average return of $e^{\mu_{S}\Delta t}$ over a time period $\Delta t$. If the reference strategy $\pi^{\text{ref}}$, which does not depend on the time horizon $T$, suggests to buy a certain amount $H_{0}$ of stocks at $t=0$, then this leads to trading costs of $H_{0}S_{0}c_{\text{prop}}$, irrespective of the time the stocks can be held. Therefore, as soon as $T$ is so small that the expected discounted profit is smaller than the trading costs, $H_{0}S_{0}(e^{\mu_{S}T}-1)<H_{0}S_{0}c_{\text{prop}}$, i.e., when $e^{\mu_{S}T}-1<c_{\text{prop}}$, the investment into the risky asset has a negative expected profit incorporating trading costs. While $\pi^{\text{ref}}$ does not take this into account, the NN strategy $\pi^{\text{NN}}_{\text{non rob}}$ learns when it is beneficial not to invest into the risky asset at all, due to this trading cost effect. In particular, we see in experiments with small $T$ that $\pi^{\text{NN}}_{\text{non rob}}\approx 0$ (Figure 8 left) [^4] Here, $e^{\mu_{S}T}-1\approx 0.0045<c_{\text{prop}}=0.01$., while it becomes more similar to $\pi^{\text{ref}}$ for larger $T$ (Figure 8 right)555 It is important to note that not all paths of $\pi^{\text{NN}}_{\text{non rob}}$ are as similar to $\pi^{\text{ref}}$ as the plotted one. For some paths, $\pi^{\text{NN}}_{\text{non rob}}$ lies above and for some it lies below the reference strategy. However, on average they seem to behave similarly. , [^6] In Figure 8, we used step sizes of $1/650$ and $1/130$ for $T=0.1$ and $T=2$, respectively.. For intermediate levels of $T$, as in the experiments in Figure 6, $\pi^{\text{NN}}_{\text{non rob}}$ seems to choose a type of interpolation between the behaviours.
@@
When decreasing the trading costs $c_{\text{prop}}$, we see a similar behaviour as when increasing $T$. Reducing $c_{\text{prop}}$ from $1\%$ to $0.1\%$ causes the NN strategy $\pi^{\text{NN}}_{\text{non rob}}$ to become more similar to $\pi^{\text{ref}}$ (Figure 9 left). This is expected because the expected profit incorporating trading costs increases with decreasing cost, as is the case for growing $T$. Further decreasing the costs, $\pi^{\text{NN}}_{\text{non rob}}$ starts to resemble $\pi_{\text{ntc}}^{\text{ref}}$ (particularly visible in terms of $H$), with a tendency to sell stocks towards the end of the time horizon (Figure 9 right). The opposite behaviour happens for increasing trading costs (being similar to decreasing $T$), where a critical cost level exists from which onwards it is not optimal to transact because the expected profit incorporating trading costs is negative. This was already found by Atkinson and Storey (2010) (and is also implied by the structure of the no-trade region (20) of Muhle-Karbe et al. (2017)). However, the authors did not draw the connection to the time horizon $T$, i.e., that the expected profit incorporating trading costs also depends on the maximal time to hold a stock and hence, that the time horizon influences whether or not it is optimal to trade.
@@
In Table 4, we see that the learned strategies $\pi^{\text{NN}}_{\text{non rob}}$ perform on par with $\pi^{\text{ref}}$ in terms of expected utility, while again attaining smaller $\operatorname{VaR}_{\alpha}$, in all cases for different time horizon $T$ and trading costs $c_{\text{prop}}$. In particular, in all cases, the NN learns an alternative strategy (as can also be seen in the plots) that has (at least) comparable performance to the reference strategy, but does not exhibit the behaviour of no-trade regions.
@@ 4.4.2 Comparison in Markets with Heavy Tailed Returns
In this section, we leverage the flexibility of our GAN-based algorithm to learn optimal investments in markets with heavier tails. To that end, we consider a market with risk-free asset $B_{t}=e^{rt}$ and one risky asset $S$, whose log-returns are independent, student-t distributed with $\nu$ degrees of freedom, location $\tilde{\mu}$ and scale $\tilde{\sigma}$. In such markets, no analytic solutions for optimal investments exist so far, even in non-robust settings for vanishing transaction costs. We therefore compare those robust and non-robust strategies to the asymptotically optimal reference strategy from Section 4.4.1.
@@
In our experiments of this section, we use a similar setup as in Section 4.4.1 with $p=0.5$, $c_{\text{prop}}=1\%$, $X_{0}=1$, $S_{0}=1$, $\tilde{\sigma}=35\%$, $\delta_{t}=1/65$ with $65$ steps. We set the drift and risk-free rate $(\tilde{\mu}_{S},r)=(7\%,3\%)$, hence, $\tilde{\mu}=10\%$, and consider two different regimes for the degrees of freedom $\nu\in\{3.5,20\}$. The robust training, model validation and model evaluation is analogous to the one explained in the experiment setup of Section 4.4.1.
@@
$M_{\tilde{u}_{p}}(\pi)$ $E_{\tilde{u}_{p}}(\pi,\tilde{\mu},\tilde{\sigma})$ $\operatorname{VaR}_{\alpha}$ $E_{\tilde{u}_{p}}(\pi,\tilde{\mu},\tilde{\sigma})$ $\operatorname{VaR}_{\alpha}$
@@
$\pi^{\text{NN}}_{\text{fully rob}}$ 2.0385 $2.0515$ $0.22$$\infty$- ∞ -
@@ Results
In Table 5, we show a comparison of the reference strategy $\pi^{\text{ref}}$ and the GAN-based strategies $\pi^{\text{NN}}$ evaluated against the reference markets (with highest, i.e., best values per column in bold). For the fully robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$, we show the results of the best combination of scaling factors, given by $\lambda_{1}=1,\lambda_{2}=10$ (for $\nu=20$).
@@
The market with $\nu=20$ degrees of freedom closely resembles the corresponding Black–Scholes market from Section 4.4.1. Accordingly, the GAN-based strategies in that market perform similarly as before: $\pi^{\text{NN}}_{\text{non rob}}$ performs comparably to $\pi^{\text{ref}}$ in terms of expected utilities $E_{\tilde{u}_{p}}$, with the additional benefit of reducing risk (reduced $\operatorname{VaR}_{\alpha}$). For GAN-based, fully robust strategies $\pi^{\text{NN}}_{\text{fully rob}}$, this effect intensifies, yielding strategies that reduce $\operatorname{VaR}_{\alpha}$ even further, at the expense of lesser expected utility $E_{\tilde{u}_{p}}$ on the reference market. However, when evaluated against the noisy market with $M_{\tilde{u}_{p}}$, $\pi^{\text{NN}}_{\text{fully rob}}$ outperforms all other strategies. Moreover, all strategies outperform the pure bank account investment $\pi^{\text{cash}}$ in $E_{\tilde{u}_{p}}$.
@@
In the market with $\nu=3.5$ degrees of freedom, we observe that the asymptotic strategy $\pi^{\text{ref}}$ leads to defaults (i.e., negative portfolio value) at some time during trading in the evaluation. The loss incurred at default leads to negative terminal portfolio values, whose utility $\tilde{u}_{p}$ is not well defined. In Table 5, we indicate negative infinity for values of $E_{\tilde{u}_{p}}$ for those strategies that can lead to defaults with non-zero probability (i.e., on at least one sample path during evaluation). Note that also with the fully robust GAN-based strategy $\pi^{\text{NN}}_{\text{fully rob}}$, we incur defaults at some point during its training. We suspect, however, that with more careful hyper-parameter selection in the GAN training, the robust strategies would not face this issue.
@@
While $\pi^{\text{ref}}$ and $\pi^{\text{NN}}_{\text{fully rob}}$ (without hyper-parameter optimization for the novel market setting) struggle to avert defaults in all market scenarios, the non-robustly trained $\pi^{\text{NN}}_{\text{non rob}}$ manages to trade without default on all scenarios even in the heavy tailed market with $\nu=3.5$ degrees of freedom. Moreover, we see in Table 5 that $\pi^{\text{NN}}_{\text{non rob}}$ outperforms the pure bank account strategy in terms of $E_{\tilde{u}_{p}}$, while keeping the VaR at a low level of $26\%$. This level is comparable to the one of the fully robust strategy in the corresponding Gaussian market (cp. $\operatorname{VaR}_{\alpha}(\pi^{\text{NN}}_{\text{fully rob}})=0.25$ in Table 3).
@@
In Figure 10, we show paths of the non-robustly trained GAN-based strategy in comparison to the asymptotically optimal solution of Section 4.4.1, on one trajectory of the heavy tailed risky asset with $\nu=20$ (left) and $\nu=3.5$ (right) degrees of freedom, respectively. While the absolute strategies (bottom plots) develop similarly in both markets, with gradually increasing numbers of risky assets over time, the GAN-based strategy learned to adapt to the market environment by investing about half as much in the risky asset in the market with $\nu=3.5$ as in the market with $\nu=20$. This highlights one of the main strongholds of our GAN-based algorithm: its robustness towards changing market environments and its flexibility to yield optimized investments even under these changing conditions.
@@ 5 Conclusion
For problems with instantaneous penalties on drift and diffusion characteristics, we substantially extended on prior work by Guo et al. (2022), by suggesting a slightly altered algorithm and testing it in a series of experiments. We showed that our GAN-based algorithm recovers known analytic solutions in friction-less market settings. Moreover, we found that with trading costs, GAN-based robust strategies outperform analytic strategies that were obtained as solutions in the friction-less setting. Therefore, GAN-based robust strategies are a valuable alternative to closed form solutions obtained in idealized settings.
@@
Overall with this paper, we found that while realistic robust utility optimization is a challenging task, GAN-based algorithms offer feasible, generic approaches to approximate otherwise unknown solutions. Going forward, it will be particularly interesting to further delve into the optimization with realistic, path-dependent penalties and employ our GAN-based algorithm to uncover potential characteristics of unknown, analytic solutions.
@@ Appendix A Additional Results for Experiments on Instantaneous Penalizations
In this section, we report on additional experiments on penalized robust utility optimization with penalties as in Equation 5. Penalties of this type tie instantaneous drift and diffusion characteristics $\mu_{t}$ and $\sigma_{t}$ to constant reference values $\tilde{\mu}$ and $\tilde{\sigma}$, respectively. A first GAN-based approach for penalized robust utility optimization of this type has been introduced in Guo et al. (2022). We extended their first experimental analyses with experiments on multi-dimensional markets, full robustification and incorporation of transaction costs in Sections 4.1 and 4.2. In the following, we report on further results.
@@ A.1 Additional Evaluation Metrics and Benchmark Strategies
As in Section 4.1, we consider the relative error $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})$ that indicates how well the penalized robust utility optimization problem is solved by a learned strategy $\pi^{\text{NN}}$. However, the relative error does not tell how well the investment strategy would work in practice, where the market usually experiences some randomness around the reference value $\tilde{\Sigma}=\tilde{\sigma}\tilde{\sigma}^{\top}$.
In Section 4.3.3, we introduced a number of approaches for modeling noisy market models for purposes of model evaluation and model selection. In this section, we revisit the “constant noise” approach of Section 4.3.3. This approach extends the idea of Guo et al. (2022) to consider a noisy (to the trader unknown) volatility matrix $\sigma^{noisy}=\tilde{\sigma}+\epsilon\tilde{\sigma}_{noise}$, where the noise has i.i.d. entries $(\tilde{\sigma}_{noise})_{i,j}\sim N(0,1)$ and $\epsilon>0$ is the noise scaling parameter. The corresponding covariance matrix is $\Sigma^{noisy}:=\sigma^{noisy}(\sigma^{noisy})^{\top}$. Similarly, we introduce a noisy drift vector $\mu^{noisy}=\tilde{\mu}+\epsilon\tilde{\mu}_{noise}$, where the noise has i.i.d. entries $(\tilde{\mu}_{noise})_{i,j}\sim N(0,1)$. We consider $\mu^{noisy}$ whenever the drift is also robustified (otherwise we use the fixed drift $\mu^{noisy}=\mu$).
As in Section 4, we compute expected utilities, now based on these noisy market parameters
@@
for the learned strategy $\pi=\pi^{\text{NN}}$ and for $\pi=\pi^{*}$. Note that in comparison to Guo et al. (2022), we do not only use one random sample for the noisy market parameters, but generate $1000$ i.i.d. samples. Then, we compute the minimum over all samples (which approximates $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}\left[u(X_{T}^{\pi})\right]$) as well as the mean, median, standard deviation and the $95\%$ approximate normal confidence interval. We compare those metrics for $\pi=\pi^{\text{NN}}$, with the ones for the analytic solution $\pi=\pi^{*}$.
@@
$\sup_{\pi}\mathbb{E}\left[u(X_{T}^{\pi,\tilde{\mu},\tilde{\Sigma}})\right],$ (21)

### v2

@@ Results
NN based strategies are able to adapt to concrete real-world settings and learn to invest more in risky assets, only if drift and diffusion characteristics allow for more potential upside over the dedicated trading time horizon, compared to the risk-less bank account investment. Depending on the time horizon, $\pi^{\text{NN}}_{\text{non rob}}$ might look very different even if annual drift and diffusion characteristics are held at the same level. The asymptotic strategy $\pi_{\text{ref}}$ however, keeps proportional investments in the no-trade region corresponding to those annual characteristics, irrespective of the time horizon. For reasonable configurations of market characteristics, the difference between the NN-based $\pi^{\text{NN}}_{\text{non rob}}$ and the asymptotic solution $\pi^{\text{ref}}$ is therefore most prominent in short-term trading.
More concretely, the risky asset $S$ has an average return of $e^{\mu_{S}\Delta t}$ over a time period $\Delta t$. If the reference strategy $\pi^{\text{ref}}$, which does not depend on the time horizon $T$, suggests to buy a certain amount $H_{0}$ of stocks at $t=0$, then this leads to trading costs of $H_{0}S_{0}c_{\text{prop}}$, irrespective of the time the stocks can be held. Therefore, as soon as $T$ is so small that the expected discounted profit is smaller than the trading costs, $H_{0}S_{0}(e^{\mu_{S}T}-1)<H_{0}S_{0}c_{\text{prop}}$, i.e., when $e^{\mu_{S}T}-1<c_{\text{prop}}$, the investment into the risky asset has a negative expected profit incorporating trading costs. While $\pi^{\text{ref}}$ does not take this into account, the NN strategy $\pi^{\text{NN}}_{\text{non rob}}$ learns when it is beneficial not to invest into the risky asset at all, due to this trading cost effect. In particular, we see in experiments with small $T$ that $\pi^{\text{NN}}_{\text{non rob}}\approx 0$ (Figure˜8 left) [^6] Here, $e^{\mu_{S}T}-1\approx 0.0045<c_{\text{prop}}=0.01$., while it becomes more similar to $\pi^{\text{ref}}$ for larger $T$ (Figure˜8 right) [^7] It is important to note that not all paths of $\pi^{\text{NN}}_{\text{non rob}}$ are as similar to $\pi^{\text{ref}}$ as the plotted one. For some paths, $\pi^{\text{NN}}_{\text{non rob}}$ lies above and for some it lies below the reference strategy. However, on average they seem to behave similarly. , [^8] In Figure 8, we used step sizes of $1/650$ and $1/130$ for $T=0.1$ and $T=2$, respectively.. For intermediate levels of $T$, as in the experiments in Figure˜6, $\pi^{\text{NN}}_{\text{non rob}}$ seems to choose a type of interpolation between the behaviours.
@@
When decreasing the trading costs $c_{\text{prop}}$, we see a similar behaviour as when increasing $T$. Reducing $c_{\text{prop}}$ from $1\%$ to $0.1\%$ causes the NN strategy $\pi^{\text{NN}}_{\text{non rob}}$ to become more similar to $\pi^{\text{ref}}$ (Figure˜9 left). This is expected because the expected profit incorporating trading costs increases with decreasing cost, as is the case for growing $T$. Further decreasing the costs, $\pi^{\text{NN}}_{\text{non rob}}$ starts to resemble $\pi_{\text{ntc}}^{\text{ref}}$ (particularly visible in terms of $H$), with a tendency to sell stocks towards the end of the time horizon (Figure˜9 right). The opposite behaviour happens for increasing trading costs (being similar to decreasing $T$), where a critical cost level exists from which onwards it is not optimal to transact because the expected profit incorporating trading costs is negative. This was already found by Atkinson and Storey (2010) (and is also implied by the structure of the no-trade region (22) of Muhle-Karbe et al. (2017)). However, the authors did not draw the connection to the time horizon $T$, i.e., that the expected profit incorporating trading costs also depends on the maximal time to hold a stock and hence, that the time horizon influences whether or not it is optimal to trade.
@@
In Table˜4, we see that the learned strategies $\pi^{\text{NN}}_{\text{non rob}}$ perform on par with $\pi^{\text{ref}}$ in terms of expected utility, while again attaining smaller $\operatorname{VaR}_{\alpha}$, in all cases for different time horizon $T$ and trading costs $c_{\text{prop}}$. In particular, in all cases, the NN learns an alternative strategy (as can also be seen in the plots) that has (at least) comparable performance to the reference strategy, but does not exhibit the behaviour of no-trade regions.
@@ 4.4.2 Comparison in Markets with Heavy Tailed Returns
In this section, we leverage the flexibility of our GAN-based algorithm to learn optimal investments in markets with heavier tails. To that end, we consider a market with risk-free asset $B_{t}=e^{rt}$ and one risky asset $S$, whose log-returns are independent, student-t distributed with $\nu$ degrees of freedom, location $\tilde{\mu}$ and scale $\tilde{\sigma}$. In such markets, no explicit solutions for optimal investments exist so far, even in non-robust settings for vanishing transaction costs. We therefore compare those robust and non-robust strategies to the asymptotically optimal reference strategy from Section˜4.4.1.
@@
In our experiments of this section, we use a similar setup as in Section˜4.4.1 with $p=0.5$, $c_{\text{prop}}=1\%$, $X_{0}=1$, $S_{0}=1$, $\tilde{\sigma}=35\%$, $\delta_{t}=1/65$ with $65$ steps. We set the drift and risk-free rate $(\tilde{\mu}_{S},r)=(7\%,3\%)$, hence, $\tilde{\mu}=10\%$, and consider two different regimes for the degrees of freedom $\nu\in\{3.5,20\}$. The robust training, model validation and model evaluation is analogous to the one explained in the experiment setup of Section˜4.4.1.
@@
[…]
@@
$\pi^{\text{NN}}_{\text{fully rob}}$ 2.0385 $2.0515$ $0.22$ -$\infty$ -
@@ Results
In Table˜5, we show a comparison of the reference strategy $\pi^{\text{ref}}$ and the GAN-based strategies $\pi^{\text{NN}}$ evaluated against the reference markets (with highest, i.e., best values per column in bold). For the fully robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$, we show the results of the best combination of scaling factors, given by $\lambda_{1}=1,\lambda_{2}=10$ (for $\nu=20$).
@@
The market with $\nu=20$ degrees of freedom closely resembles the corresponding Black–Scholes market from Section˜4.4.1. Accordingly, the GAN-based strategies in that market perform similarly as before: $\pi^{\text{NN}}_{\text{non rob}}$ performs comparably to $\pi^{\text{ref}}$ in terms of expected utilities $E_{\tilde{u}_{p}}$, with the additional benefit of reducing risk (reduced $\operatorname{VaR}_{\alpha}$). For GAN-based, fully robust strategies $\pi^{\text{NN}}_{\text{fully rob}}$, this effect intensifies, yielding strategies that reduce $\operatorname{VaR}_{\alpha}$ even further, at the expense of lesser expected utility $E_{\tilde{u}_{p}}$ on the reference market. However, when evaluated against the noisy market with $M_{\tilde{u}_{p}}$, $\pi^{\text{NN}}_{\text{fully rob}}$ outperforms all other strategies. Moreover, all strategies outperform the pure bank account investment $\pi^{\text{cash}}$ in $E_{\tilde{u}_{p}}$.
@@
In the market with $\nu=3.5$ degrees of freedom, we observe that the asymptotic strategy $\pi^{\text{ref}}$ leads to defaults (i.e., negative portfolio value) at some time during trading in the evaluation. The loss incurred at default leads to negative terminal portfolio values, whose utility $\tilde{u}_{p}$ is not well defined. In Table˜5, we indicate negative infinity for values of $E_{\tilde{u}_{p}}$ for those strategies that can lead to defaults with non-zero probability (i.e., on at least one sample path during evaluation). Note that also with the fully robust GAN-based strategy $\pi^{\text{NN}}_{\text{fully rob}}$, we incur defaults at some point during its training. We suspect, however, that with more careful hyper-parameter selection in the GAN training, the robust strategies would not face this issue.
@@
While $\pi^{\text{ref}}$ and $\pi^{\text{NN}}_{\text{fully rob}}$ (without hyper-parameter optimization for the novel market setting) struggle to avert defaults in all market scenarios, the non-robustly trained $\pi^{\text{NN}}_{\text{non rob}}$ manages to trade without default on all scenarios even in the heavy tailed market with $\nu=3.5$ degrees of freedom. Moreover, we see in Table˜5 that $\pi^{\text{NN}}_{\text{non rob}}$ outperforms the pure bank account strategy in terms of $E_{\tilde{u}_{p}}$, while keeping the VaR at a low level of $26\%$. This level is comparable to the one of the fully robust strategy in the corresponding Gaussian market (cp. $\operatorname{VaR}_{\alpha}(\pi^{\text{NN}}_{\text{fully rob}})=0.25$ in Table˜3).
@@
In Figure˜10, we show paths of the non-robustly trained GAN-based strategy in comparison to the asymptotically optimal solution of Section˜4.4.1, on one trajectory of the heavy tailed risky asset with $\nu=20$ (left) and $\nu=3.5$ (right) degrees of freedom, respectively. While the absolute strategies (bottom plots) develop similarly in both markets, with gradually increasing numbers of risky assets over time, the GAN-based strategy learned to adapt to the market environment by investing about half as much in the risky asset in the market with $\nu=3.5$ as in the market with $\nu=20$. This highlights one of the main strongholds of our GAN-based algorithm: its robustness towards changing market environments and its flexibility to yield optimized investments even under these changing conditions.
5 A Note on the Path-Dependence of the Trading and Market Strategy
In this section, we discuss in more detail the path-dependence of the optimal trading strategy and market characteristics in the context of penalized robust utility optimization with path-dependent penalties and transaction costs. Trading costs and path-dependent penalties are the two factors increasing the complexity of the robust utility optimization problem such that a (Markovian) explicit solution as in Section˜4.1 can no longer be expected.
5.1 Path-Dependent Penalties
[…]
5.2 Transaction Costs
The asymptotically optimal trading strategy for small transaction costs, which was discussed in Section˜4.4, has a generic path-dependence in terms of the inputs that we use. In particular, it needs knowledge of the amount of stocks $H_{t_{n-1}}$ held at the previous time step $t_{n-1}$ to decide whether to keep the same number of stocks or to move with the boundary of the no-trade region. While the strategy could be made Markovian by including $H$ as an input (state) variable for the strategy, our setting only uses $S,X$ as such. Clearly, $H$ can be reconstructed from the history of the path of $S$ and $X$ such that the path-dependent RNN strategy could in principle learn the asymptotically optimal strategy, while this is impossible for the Markovian FFNN strategy. Nevertheless, we saw that the asymptotically optimal strategy is not always optimal in settings with small transaction costs, in particular when the time horizon is small. In this case, the optimal strategy is Markovian and simply keeps everything in the risk free asset. In cases where the time horizon is larger, it is not clear what the true optimal strategy looks like. Our results showed that the learned Markovian strategies achieve the same results in terms of expected utility, while using a qualitatively different strategy. This suggests the hypothesis that Markovian strategies are dense in the set of all strategies, where the pseudo-distance between strategies is measure in terms of difference in expected utility.
@@ 6 Conclusion
For problems with instantaneous penalties on drift and diffusion characteristics, we substantially extended on prior work by Guo et al. (2022), by suggesting a slightly altered algorithm and testing it in a series of experiments. We showed that our GAN-based algorithm recovers known explicit solutions in friction-less market settings. Moreover, we found that with trading costs, GAN-based robust strategies outperform explicit strategies that were obtained as solutions in the friction-less setting. Therefore, GAN-based robust strategies are a valuable alternative to closed form solutions obtained in idealized settings.
@@
Overall with this paper, we found that while realistic robust utility optimization is a challenging task, GAN-based algorithms offer feasible, generic approaches to approximate otherwise unknown solutions. Going forward, it will be particularly interesting to further delve into the optimization with realistic, path-dependent penalties and employ our GAN-based algorithm to uncover potential characteristics of unknown, explicit solutions.
Declaration of funding
No funding was received.
Disclosure of interest
There are no potential conflicts of interest.
@@ Appendix A Additional Results for Experiments on Instantaneous Penalizations
In this section, we report on additional experiments on penalized robust utility optimization with penalties as in Equation˜5. Penalties of this type tie instantaneous drift and diffusion characteristics $\mu_{t}$ and $\sigma_{t}$ to constant reference values $\tilde{\mu}$ and $\tilde{\sigma}$, respectively. A first GAN-based approach for penalized robust utility optimization of this type has been introduced in Guo et al. (2022). We extended their first experimental analyses with experiments on multi-dimensional markets, full robustification and incorporation of transaction costs in Sections˜4.1 and 4.2. In the following, we report on further results.
@@ A.1 Additional Evaluation Metrics and Benchmark Strategies
As in Section˜4.1, we consider the relative error $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})$ that indicates how well the penalized robust utility optimization problem is solved by a learned strategy $\pi^{\text{NN}}$. However, the relative error does not tell how well the investment strategy would work in practice, where the market usually experiences some randomness around the reference value $\tilde{\Sigma}=\tilde{\sigma}\tilde{\sigma}^{\top}$.
In Section˜4.3.3, we introduced a number of approaches for modeling noisy market models for purposes of model evaluation and model selection. In this section, we revisit the “constant noise” approach of Section˜4.3.3. This approach extends the idea of Guo et al. (2022) to consider a noisy (to the trader unknown) volatility matrix $\sigma^{noisy}=\tilde{\sigma}+\epsilon\tilde{\sigma}_{noise}$, where the noise has i.i.d. entries $(\tilde{\sigma}_{noise})_{i,j}\sim N(0,1)$ and $\epsilon>0$ is the noise scaling parameter. The corresponding covariance matrix is $\Sigma^{noisy}:=\sigma^{noisy}(\sigma^{noisy})^{\top}$. Similarly, we introduce a noisy drift vector $\mu^{noisy}=\tilde{\mu}+\epsilon\tilde{\mu}_{noise}$, where the noise has i.i.d. entries $(\tilde{\mu}_{noise})_{i,j}\sim N(0,1)$. We consider $\mu^{noisy}$ whenever the drift is also robustified (otherwise we use the fixed drift $\mu^{noisy}=\mu$).
As in Section˜4, we compute expected utilities, now based on these noisy market parameters
@@
for the learned strategy $\pi=\pi^{\text{NN}}$ and for $\pi=\pi^{*}$. Note that in comparison to Guo et al. (2022), we do not only use one random sample for the noisy market parameters, but generate $1000$ i.i.d. samples. Then, we compute the minimum over all samples (which approximates $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}\left[u(X_{T}^{\pi})\right]$) as well as the mean, median, standard deviation and the $95\%$ approximate normal confidence interval. We compare those metrics for $\pi=\pi^{\text{NN}}$, with the ones for the explicit solution $\pi=\pi^{*}$.
@@
[…]